Order posts by limited to posts

24 Feb 2014 16:00:00
24 Feb 2014 16:01:15
Some work was done today (following issues at the weekend). This work is on part of a redundant pair of routers, and so should have no impact. The normal redundancy aspects have worked, with changes to routing and VRRP, but for some unknown reason there are issues with BGP announcements internally which are causing blips on external connectivity to some Ethernet customers. The work has been completed but we are now investigating how this has had an impact on services so they can be avoided in future.
Started 24 Feb 2014 15:45:00
Closed 24 Feb 2014 16:00:00
Previously expected 24 Feb 2014 16:00:00

02 Sep 2011 09:41:55
02 Sep 2011 09:43:14

The usage graphs were disrupted yesterday due to various work on core routers, and will be reset today starting shortly. This is to fix a problem where the ping monitoring was not working correctly. Service should otherwise be unaffected by this change.

Started 02 Sep 2011
Closed 02 Sep 2011 09:41:55